# Exploiting Meta-Learning-based Poisoning Attacks for Graph Link Prediction

### v1

@@ Abstract
Link prediction in graph data utilizes various algorithms and machine learning/deep learning models to predict potential relationships between graph nodes. This technique has found widespread use in numerous real-world applications, including recommendation systems, community networks, and biological structures. However, recent research has highlighted the vulnerability of link prediction models to adversarial attacks, such as poisoning and evasion attacks. Addressing the vulnerability of these models is crucial to ensure stable and robust performance in link prediction applications. While many works have focused on enhancing the robustness of the Graph Convolution Network (GCN) model, the Variational Graph Auto-Encoder (VGAE), a sophisticated model for link prediction, has not been thoroughly investigated in the context of graph adversarial attacks. To bridge this gap, this article proposes an unweighted graph poisoning attack approach using meta-learning techniques to undermine VGAE’s link prediction performance. We conducted comprehensive experiments on diverse datasets to evaluate the proposed method and its parameters, comparing it with existing approaches in similar settings. Our results demonstrate that our approach significantly diminishes link prediction performance and outperforms other state-of-the-art methods.
{IEEEImpStatement}
Link prediction is a popular and critical technology in many applications, it helps organizations and researchers to understand and analyze the graph network. However, recent researches have pointed out that the link prediction model is vulnerable under adversarial attacks. Addressing the vulnerability of link prediction model can enhance the robustness of the link prediction model. The adversarial attack proposed in this paper reveals the vulnerabilities of a popular link prediction model (VGAE). With a minor and unnoticeable modification on the graph structure, the proposed adversarial approach can successfully degrade the performance of link prediction model. This approach indicates the possible threat to link prediction applications. It could provide a direction for enhancing the robustness of link prediction models.
{IEEEkeywords}
Adversarial attack, Link prediction, Meta-learning, Variational Graph Auto-encoder (VGAE)
1 Introduction
@@
Many real-world datasets can be represented as graphs containing nodes, links, and associated features. Examples include social networks [1], traffic networks [2], biological structures [3], e-commerce networks [4], and so on. The structure of these networks is dynamic, changing over time as relationships evolve. Therefore, capturing these changes in the underlying graph is crucial for any application. Link prediction addresses this challenge and finds applications in various real-world scenarios, such as recommendation systems [5]. It employs diverse algorithms and machine learning (ML)/deep learning (DL) models to predict potential relationships (links) between nodes in a graph.
Over the past few decades, numerous approaches to link prediction have emerged across various research domains, including Local Random Walk [6], DeepWalk [7], node2vec [8], Graph Convolutional Networks (GCN) [9], Variational Graph Auto-encoder (VGAE) [10], and others. However, recent studies have underscored the susceptibility of graph learning models to adversarial attacks [11], such as poisoning and evasion attacks, which pose significant threats to the performance and stability of link prediction models. Given the widespread use of link prediction in real-world applications, ensuring the robustness and stability of these algorithms is essential. In essence, addressing the vulnerability of link prediction models is crucial, and this often falls within the realm of graph adversarial attacks.
@@
In recent years, numerous approaches to graph adversarial attacks have been proposed. Xu et al. [12] proposed a projected gradient descent (PGD) topology attack from an optimization perspective to optimize the negative cross-entropy using the gradient to maximize the prediction error of the model. Waniek et al. [13] proposed the Disconnect Internally, Connect Externally (DICE) algorithm, which reduces internal graph density to deceive graph models. Chen et al. [14] leveraged an adversarial network generator to initiate attacks based on gradients, effectively misleading state-of-the-art link prediction algorithms. Furthermore, Meta-learning [15], offering valuable insights for efficient learning and rapid adaptation, has emerged as an attack strategy. In [16], the authors utilized meta-learning to conduct adversarial modifications on graph data.
Although minor alterations to graph data can significantly affect model performance, crafting an effective graph adversarial attack remains challenging. Attackers must meticulously design their strategies to ensure effectiveness. Additionally, as models and learning techniques advance, modern deep learning models exhibit increasing robustness, surpassing the resilience of previous generation machine learning models. For example, in [14], the author highlights the superior robustness of Graph Auto Encoder (GAE) compared to other state-of-the-art link prediction models. In addition, Variational Graph Auto-encoder (VGAE), which can be considered as an advancement of GAE has shown competitive performance on link prediction tasks. It uses a probabilistic approach, allowing the model to learn more flexible and interpretable latent representations of graph data. Nevertheless, the adversarial attack on VGAE remains a significant and open challenge.
@@
In this study, we aim to bridge this gap by investigating adversarial attacks on VGAE and propose a solution through the introduction of an unweighted graph poisoning attack approach. Leveraging meta-learning techniques during the adversarial training stage, our method aims to degrade the link prediction performance of VGAE. Through extensive experiments on diverse datasets, we evaluated our approach and its parameters. In addition, we compare our method with existing approaches, demonstrating its superior effectiveness. To ensure reproducibility and facilitate further research, we have made the source code publicly available at our github repository [^1] https://github.com/mingchenli/VGAE_Attack_Meta.
The main contributions of our approach are summarized as follows:
@@
  • We introduce an innovative poisoning attack leveraging meta-learning to effectively degrade the link prediction performance of VGAE.
  • By leveraging meta-learning, we determine optimal attack directions resulting in subtle modifications in the graph, while defining the adversarial behavior of these modifications.
  • Extensive experiments were conducted to assess the impact of parameters and the effectiveness of our approach across three benchmark datasets under various scenarios. Furthermore, we compare our approach with the state-of-the-art approaches to gauge its impact.
@@
The remaining sections of the paper are structured as follows: Section 2 provides an overview of related works, Section 3 offers background knowledge on our approach, Section 4 outlines our methodology, and Section 5 presents the experimental evaluation, and Section 6 concludes the paper.
2 Related Works
@@
2.1 Gradient-Based Attacks
Gradient-based adversarial attacks utilize model gradients to identify and perturb critical graph elements. Xu et al. [12] proposed projected gradient descent (PGD) and MinMax attacks that treat the attack as an optimization problem. They have proposed an optimization framework to optimize the negative cross-entropy using the gradient to maximize the prediction error of the model. Their method uses projected gradient descent (PGD) to iteratively modify links, significantly affecting model performance while respecting a budget constraint on the number of modifications. Chen et al. [14] introduced the iterative gradient attack (IGA) for the attack on link prediction tasks. This method leverages gradients derived from trained graph auto-encoders to identify and perturb the significant links in the graph. They perform comprehensive evaluations to reveal that even minor link alterations could disrupt the performance of state-of-the-art link prediction models. Zhang et al. [17] developed a gradient-based attack based on graph contrastive learning. By maximizing the contrastive loss during the training phase, the attack disrupts the embeddings produced by the contrastive model under an unsupervised learning setting, making it suitable for real-world where sometimes the data labels are scarce. In [16], a meta-learning based poisoning attack strategy was proposed to target the significant part of the graph structure or node characteristics. They using meta-gradient as guidance to generate adversarial modifications to decrease graph model performance.
2.2 Heuristic-Based Attacks
Heuristic-based attacks rely on predefined strategies to perturb graphs, often utilizing insights from graph theory. Zügner et al. proposed Nettack  [18], a framework that uses a greedy algorithm to identify and perturb key links and node attributes, forcing misclassifications in node classification tasks. Nettack shows that the weaknesses of the models are highly dependent on graph structure in semi-supervised settings. DICE attack introduced in [13], uses the strategies Disconnect Internally, Connect Externally (DICE) to manipulate the internal density within communities and disrupt specific targets in the graph. The attack modifies the graph structure strategically to jeopardize the performance of the graph model.
2.3 Reinforcement Learning-Based Attacks
@@
However, most of these attacks mentioned above are targeted or designed for graph convolutional networks (GCNs) [9], which serve as the foundational model for graph data analysis. GCN has certain limitations on its representation capacity, graph data types, and generative capabilities. In contrast, VGAE has become a popular and more powerful model than GCNs for the link prediction task. Our approach presents a distinct meta-learning framework and proposes predefined attack modifications tailored for adversarial attacks in link prediction.
3 Background
@@
3.1 Graph Adversarial Attack
In general, graph adversarial attacks utilize intentional modifications or perturbations on the graph structure, node attributes, or relevant features to degrade the performance of the target model. For graph data, these common modifications can be applied to the node, links (edges), and the whole graph, such as adding or deleting nodes of links, changing link weights (in weighted graph), modifying node features, and generating fake graphs.
@@ 3.1 Graph Adversarial Attack
Graph adversarial attacks can target machine learning/deep learning models at various stages throughout the learning pipeline. Based on the target stage, these attacks can be broadly categorized into two main categories (as shown in Fig. 1):
@@
In this paper, our goal is to design a poisoning attack on link prediction tasks for both untargeted and target scenarios.
3.2 Variational Graph Auto-Encoders (VGAE)
@@
In summary, the Variational Graph Autoencoder (VGAE) is a powerful generative model that learns to encode data into a probabilistic latent space and decode from this space to reconstruct the graph to perform link prediction or node classification tasks. Using KL divergence as a regularization term aims to balance reconstruction accuracy while ensuring that the latent space has known distribution and meaningful sampling. Due to these characteristics, the VGAE often outperforms conventional GCN-based approaches for link prediction, as it better captures the underlying distribution of the graph data. Hence, it is worth investigating the adversarial attack on VGAE.
3.3 Adversarial Attack with Meta-Learning
@@
The concept of meta-learning can also be applied to graph adversarial attacks. As discussed in Section 2, attackers can leverage the meta-gradient obtained during training towards their target [16]. The meta-gradient is computed by unrolling the training process into multiple tasks, which can help the attack learn a more holistic picture of the model than looking into any certain phase in the training. By leveraging this meta-gradient, attackers gain a powerful road map, enabling them to manipulate graph data effectively and apply a powerful attack strategy. More specifically, the attacker can identify the node or links with the greatest impact during model learning and perform the precise attack instead of making random or naive perturbations on the graph. In this paper, we will take the advantage of the meta-learning to design our poisoning attack strategies.
@@ 4 Methodology
4.1 Objective
In the proposed adversarial attack approach, the attacker utilizes meta-learning during the training phase of the attack model to make subtle modifications to the unweighted graph structure (adjacency matrix, denoted as $Adj$). As depicted in Fig. 2, these modifications generate poisoned data for the poisoning attack, which involves training the target model with the poisoned data—a process known as poisoning training.
The goal of this approach is to diminish the prediction accuracy of the model specifically on the targeted link(s). In other words, our objective is to reduce the link prediction accuracy of specified link(s) on Variational Graph Auto-Encoders through the poisoning attack.
4.2 Attack Setting and Threat Model
4.2.1 Attacker’s Goal
The attacker aims to increase the misclassification rate of the prediction of VGAE on a specific link (s) using the proposed poisoning attack. By introducing malicious and inconspicuous changes to the graph data, the attacker tries to make the learned VGAE model inherently biased after training on the poisoning data. For example, the goal is to induce the model to predict a link between two unrelated nodes, which should not be predicted.
4.2.2 Attacker’s Knowledge
@@
  • Model Knowledge: The attacker does not know the exact model architecture, such as the number of layers and hyperparameters of the model. They only know that the VGAE or GNN-based model is used for link prediction.
  • Training Pipeline Knowledge: The attack has a general understanding of when the model is trained or updated to time the poisoning attack. However, they do not have control or alter the original training procedure (i.e., the optimizer, learning rate, or objective function).
4.2.3 Attacker’s Capacity
@@
4.2.4 Threat Model
Leveraging the attacker’s knowledge and capacity, the attacker selectively adds or removes links within the allowed budget to generate poisoning data via their own framework (Section 4.3). These poisoned data distort the structural patterns of the graph, causing the VGAE to learn misleading latent representations. In other words, the attacker relies on subtle but carefully chosen structural modification on the graph data to achieve their goal of inducing incorrect link prediction.
4.3 Attack Framework
Fig. 2 provides an overview of our approach. It begins with the original graph data (original $Adj$ and original features) as input and generates poisoned graph data (poisoned $Adj$ and original features) for poisoning training. Our poisoning attack consists of two primary components that operate sequentially to produce the poisoned graph data for conducting poisoning training on VGAE:
  • Attack model: The attacker employs a surrogate model as an attack model to train on the original graph data. During the training, meta-learning with meta-gradient is utilized by the attacker to identify effective modifications to the graph structure.
  • Graph structure modifier: Following the training of the attack using meta-learning, the attacker employs a graph structure modifier to adjust the original adjacency matrix (original $Adj$) and create an adversarial adjacency matrix (adversarial $Adj$). Subsequently, the attacker combines the adversarial $Adj.$ with the original graph features to generate poisoned graph data for poisoning training.
Input: Attack budget $\Delta$, graph data $G(A,X)$, number of epoches $k$, attack model $M$

### v2

@@ Abstract
Link prediction in graph data uses various algorithms and Graph Nerual Network (GNN) models to predict potential relationships between graph nodes. These techniques have found widespread use in numerous real-world applications, including recommendation systems, community/social networks, and biological structures. However, recent research has highlighted the vulnerability of GNN models to adversarial attacks, such as poisoning and evasion attacks. Addressing the vulnerability of GNN models is crucial to ensure stable and robust performance in GNN applications. Although many works have focused on enhancing the robustness of node classification on GNN models, the robustness of link prediction has received less attention. To bridge this gap, this article introduces an unweighted graph poisoning attack that leverages meta-learning with weighted scheme strategies to degrade the link prediction performance of GNNs. We conducted comprehensive experiments on diverse datasets across multiple link prediction applications to evaluate the proposed method and its parameters, comparing it with existing approaches under similar conditions. Our results demonstrate that our approach significantly reduces link prediction performance and consistently outperforms other state-of-the-art baselines.
I Introduction
@@
Many real-world datasets such as social networks [1], traffic networks [2], biological structures [3], and e-commerce networks [4] can be represented as graphs containing nodes, links, and associated features. The structure of these graphs is dynamic, constantly changing as relationships evolve. Capturing these changes is fundamental to the success of various real-world scenarios, such as community detection and recommendation systems [5]. Node classification and link prediction are the primary tasks that address this challenge, employing various algorithms and graph neural network models to predict potential labels (nodes) and relationships (links) between nodes in a graph.
@@
During the past few decades, numerous graph approaches have emerged in various research domains, including Local Random Walk [6], DeepWalk [7], node2vec [8], Graph Convolutional Networks (GCN) [9], Variational Graph Autoencoder (VGAE) [10], Graph Attention Network (GAT) [11] and others. However, recent studies have underscored the susceptibility of graph learning models to adversarial attacks [12], such as poisoning and evasion attacks, which pose significant threats to the performance and stability of graph models. To exploit these vulnerabilities, numerous approaches of graph adversarial attacks have been proposed. Xu et al. [13] proposed a topology attack with projected gradient descent (PGD) from an optimization perspective to optimize the negative cross-entropy using the gradient to maximize the prediction error of the model. Waniek et al. [14] proposed the Disconnect Internally, Connect Externally (DICE) algorithm, which reduces the internal graph density to deceive graph models. Chen et al. [15] used an adversarial network generator to initiate gradient-based attacks, effectively misleading state-of-the-art link prediction algorithms. Furthermore, Meta-learning [16], offering valuable insights for efficient learning and rapid adaptation, has emerged as an attack strategy. In [17], the authors utilized meta-learning to conduct adversarial modifications on graph data. A key observation is that most existing methods primarily focus on node classification tasks, while the adversarial vulnerability of link prediction remains comparatively underexplored. This represents a significant and important gap in the current research landscape.
@@
In this study, we aim to bridge this gap by investigating adversarial attacks on graph link prediction and propose an unweighted graph poisoning attack approach. Leveraging meta-learning techniques with weighted schemes during the adversarial training stage, our method aims to degrade the performance of link prediction models. Through extensive experiments on diverse datasets and corresponding use cases, we evaluated the effectiveness of our approach and the impact of the parameters. In addition, we compare our method with existing approaches, demonstrating its superior effectiveness. To ensure reproducibility and facilitate further research, we have made the source code publicly available at our github repository [^1] https://github.com/mingchenli/VGAE_Attack_Meta.
The main contributions of this work are as follows:
@@
  • We propose a novel poisoning attack that leverages meta-learning with weighted schemes to effectively degrade link-prediction performance.
  • We introduce three weighted schemes within the meta-learning framework to identify optimal attack directions, enabling subtle yet impactful adversarial graph modifications.
  • We present a comparative analysis of these weighted schemes, experimentally validating our design and providing key insights for the future development of weighted strategies.
  • We conduct extensive experiments on benchmark datasets across diverse use cases, analyzing parameter effects and demonstrating that our method outperforms state-of-the-art approaches.
@@
The remaining sections of the paper are structured as follows: Section II provides an overview of related works, Section III offers background knowledge on our approach, Section IV outlines our methodology, Section V presents the experimental evaluation, and Section VI concludes the paper.
II Related Works
@@
II-A Gradient-Based Attacks
Gradient-based adversarial attacks utilize model gradients to identify and perturb critical graph elements. Xu et al. [13] proposed projected gradient descent (PGD) and MinMax attacks that treat the attack as an optimization problem. They have proposed an optimization framework to optimize the negative cross-entropy using the gradient to maximize the prediction error of the model. Their method uses PGD to iteratively modify links, significantly affecting model performance while respecting a budget constraint on the number of modifications. Chen et al. [15] introduced the iterative gradient attack (IGA) for the attack on link prediction tasks. This method leverages gradients derived from trained graph auto-encoders to identify and perturb the significant links in the graph. They perform comprehensive evaluations to demonstrate that even minor link alterations can disrupt the performance of state-of-the-art link prediction models. Zhang et al. [18] developed a gradient-based attack based on graph contrastive learning. By maximizing the contrastive loss during the training phase, the attack disrupts the embeddings produced by the contrastive model under an unsupervised learning setting, making it suitable for real-world where sometimes the data labels are scarce. In [17], a meta-learning based poisoning attack strategy was proposed to target the significant part of the graph structure or node characteristics. They used meta-gradient as guidance to generate adversarial modifications to decrease graph model performance.
II-B Heuristic-Based Attacks
Heuristic-based attacks employ predefined strategies to perturb graphs, often drawing on principles from graph theory. Zügner et al. proposed Nettack [19], a framework that applies a greedy algorithm to identify and modify critical links and node attributes, thereby inducing misclassifications in node-classification tasks. Their results show that model vulnerabilities are strongly tied to graph structure in semi-supervised settings.The DICE attack introduced in [14] follows the Disconnect Internally, Connect Externally (DICE) strategy to manipulate community density and disrupt specific targets. By strategically altering the graph structure, DICE effectively degrades the performance of graph-based models.
II-C Reinforcement Learning-Based Attacks
@@
However, most of the attacks described above are aimed at graph convolutional networks [9] and node classification. GCNs face limitations in representational capacity, supported graph types, and generative abilities. On the other hand, link prediction is an equally important and widely studied task in graph learning. Many models like VGAE, LightGCN and GAT have emerged as the more powerful models for link prediction. Our work introduces a distinct weighted meta-learning framework with carefully designed attack modifications tailored specifically for adversarial link prediction tasks.
III Background
@@
III-A Graph Adversarial Attack
In general, graph adversarial attacks deliberately modify the graph structure, node attributes, or related features to degrade the performance of a target model. Typical modifications include adding or deleting nodes or edges, altering link weights in weighted graphs, changing node features, or generating synthetic graphs. Such attacks can target machine-learning or deep-learning models at various stages of the learning pipeline. Based on the targeted stage, they can be broadly classified into two main categories (as illustrated in Fig. 1).
@@ III-A Graph Adversarial Attack
In this paper, our goal is to design a poisoning attack on link prediction tasks for untargeted scenarios.
III-B Graph Neural Networks
III-B1 Variational Graph Auto-Encoders (VGAE)
@@
III-B2 Graph Attention Network (GAT)
The Graph Attention Network (GAT), introduced by Veličković et al. [11], enhances graph representation learning by incorporating an attention mechanism. Unlike conventional graph convolutional networks that aggregate neighbor information uniformly, GAT allows nodes to assign different levels of importance to their neighbors. This is achieved by computing attention coefficients for each edge, which dynamically weigh the contribution of neighboring node features during the aggregation process. The core operation of a GAT layer can be expressed as:
$\vec{h^{\prime}}_{i}=\sigma\left(\sum{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{W}\vec{h}_{j}\right)$ (2)
where $\vec{h^{\prime}}_{i}$ is the updated feature vector for node $i$, $\sigma$ is a non-linear activation function, $\mathcal{N}_{i}$ represents the neighbors of node $i$, $W$ is a shared weight matrix for feature transformation, and $\alpha_{ij}$ is the normalized attention coefficient of node $j$ to node $i$. The attention coefficients are computed based on the features of the connected nodes, allowing the model to focus on the most relevant parts of the neighborhood.
III-B3 LightGCN
LightGCN, proposed by He et al. [23], is a simplified graph convolutional network designed specifically for recommendation and collaborative filtering. LightGCN simplifies the design by preserving neighborhood aggregation which is the essential part in GCN. It learns user and item embeddings by linearly propagating them on the user-item interaction graph. The aggregation at layer $k+1$ is defined as:
$\mathbf{e}_{u}^{(k+1)}=\sum{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}|}\sqrt{|\mathcal{N}_{i}|}}\mathbf{e}_{i}^{(k)}$ (3)
where $\mathbf{e}_{u}^{(k)}$ and $\mathbf{e}_{i}^{(k)}$ are the embeddings of user u and item i at layer k, respectively, and the term with square roots is a symmetric normalization factor. The final embedding for a user is a weighted sum of its embeddings from all layers, which effectively combines information from its multi-hop neighbors. This simple linear propagation captures the collaborative filtering effect smoothly.
III-C Adversarial Attack with Meta-Learning
@@
The concept of meta-learning can also be applied to graph adversarial attacks. As discussed in Section II, attackers can leverage the meta-gradient obtained during training towards their target [17]. The meta-gradient is computed by unrolling the training process into multiple tasks, which can help the attack learn a more holistic picture of the model than looking into any certain phase in the training. By leveraging this meta-gradient, attackers gain a powerful road map, enabling them to manipulate graph data effectively and apply a powerful attack strategy. More specifically, the attacker can identify the node or links with the greatest impact during model learning and perform the precise attack instead of making random or naive perturbations on the graph. In this paper, we will take advantage of meta-learning to design our poisoning attack strategies.
@@ IV Methodology
IV-A Objective
In the proposed adversarial attack approach, the attacker utilizes meta-learning with the weighted scheme during the training phase of the attack model to make subtle modifications to the unweighted graph structure (adjacency matrix, denoted as $Adj$). As depicted in Fig. 2, these modifications generate poisoned data for the poisoning attack, which involves training the victim models with the poisoned data, named poisoning training.
The goal of this approach is to reduce the performance of the GNN models. In other words, our objective is to propose a model-agnostic poisoning attack to reduce the performance of link prediction models.
IV-B Attack Setting and Threat Model
IV-B1 Attacker’s Goal
The attacker aims to increase the error rate of the link prediction models using the proposed poisoning attack. By introducing malicious and inconspicuous changes to the graph data, the attacker tries to make the learned link prediction models inherently biased after training on the poisoning data. For example, the goal is to induce the model to predict a link between two unrelated nodes, which should not be predicted.
IV-B2 Attacker’s Knowledge
@@
  • Model Knowledge: The attacker does not know the exact model architecture, such as the number of layers and hyperparameters of the model.
  • Training Pipeline Knowledge: The attack has a general understanding of when the model is trained or updated. However, they do not have control or alter the original training procedure (i.e., the optimizer, learning rate).
IV-B3 Attacker’s Capacity
@@
IV-B4 Threat Model
Leveraging the attacker’s knowledge and capacity, the attacker selectively adds or removes links within the allowed budget to generate poisoning data via their own framework (Section IV-C). These poisoned data distort the structural patterns of the graph, causing the link prediction models to learn misleading latent representations. In other words, the attacker relies on subtle but carefully chosen structural modification on the graph data to achieve their goal of inducing incorrect link prediction.
IV-C Attack Framework
Fig. 2 provides an overview of our approach. It begins with the original graph data (original $Adj$ and original features) as input and generates poisoned graph data (poisoned $Adj$ and original features) for poisoning training. Our adversarial procedure consists of two primary components that operate sequentially to produce the poisoned graph data to produce poisoned data.
  • Attack model: The attacker employs a surrogate model as an attack model to train on the original graph data. During the surrogate model training, the attacker utilizes the weighted meta-learning to construct the meta-gradient. This meta-gradient helps the attacker to identify effective modifications to the graph structure.
  • Graph structure modifier: Following the training of the attack model, the attacker employs a graph structure modifier to adjust the original adjacency matrix (original $Adj$) and create an adversarial adjacency matrix (adversarial $Adj$) based on meta-gradient. Subsequently, the attacker combines the adversarial $Adj.$ with the original graph features to generate poisoned graph data for poisoning training.
Input: Attack budget $\Delta$, graph data $G(A,X)$, number of epoches $k$, attack model $M$, weight mechanism $W$